# jina-vlm: Small Multilingual Vision Language Model

### v1

@@ Abstract
We present jina-vlm, a 2.4B parameter vision-language model that achieves state-of-the-art multilingual visual question answering among open 2B-scale VLMs. The model couples a SigLIP2 vision encoder with a Qwen3 language backbone through an attention-pooling connector that enables token-efficient processing of arbitrary-resolution images. Across standard VQA benchmarks and multilingual evaluations, jina-vlm outperforms comparable models while preserving competitive text-only performance.
@@ 1 Introduction
This work introduces jina-vlm, a 2.4B parameter VLM that addresses both challenges. The model aligns a SigLIP2-So400M/14-384 vision encoder (siglip2) with Qwen3-1.7B-Base (qwen3) through an attention-pooling connector, trained with a two-stage pipeline that explicitly incorporates multilingual data. Among open 2B-scale VLMs, jina-vlm achieves state-of-the-art performance on multilingual multimodal benchmarks including MMMB and Multilingual MMBench, demonstrating that small models can excel at cross-lingual visual understanding without sacrificing general capabilities. The model matches or exceeds comparable VLMs on standard English benchmarks spanning diagrams, charts, documents, and OCR, achieving the highest average score (72.3) across eight VQA benchmarks. These results are enabled by two technical contributions: an efficient arbitrary-resolution pipeline that combines overlapping tiling with attention-based token pooling to reduce visual token count by 4, and a training recipe that incorporates text-only data to preserve the language understanding performance of the backbone LLM.
@@ 2 Related Work
Efficient resolution-agnostic image processing. Standard ViTs process fixed-resolution images, requiring resizing that discards fine-grained detail. Since visual token count scales with resolution and Transformer computation scales quadratically with sequence length, naive high-resolution processing is prohibitive. Several solutions exist: molmo tile images with overlap; qwen_2_vl introduce Naive Dynamic Resolution with Multimodal Rotary Position Embedding (rope; 2drope); ovis_2_5 use native-resolution ViTs (navit). Orthogonally, images often contain low-information regions (e.g., sky backgrounds), making visual tokens highly redundant. Token compression methods address this (fastv; prumerge; visionzip; pyramiddrop). internvl_1_5 develop Dynamic High-Resolution Tiling, and nvila propose scale-then-compress strategies.
@@ 5.1 General VQA Tasks
Results for models other than jina-vlm are from their respective papers, internvl_3_5; internvl_3; qwen_2_vl, except those marked with * which were computed using VLMEvalKit. All scores represent accuracy (%) except OCRBench which uses a 0-1000 scale, normalized to 0-100 to align scales with the other benchmarks.
@@ 5.2 Document and Real-World Understanding
Results for models other than jina-vlm are from their respective papers, internvl_3_5; internvl_3; qwen_2_vl, except those marked with * which are computed using VLMEvalKit. All scores represent accuracy (%) except MME which uses a 0-2800 scale, normalized to 0-100 to align scales with the other benchmarks.
@@ 5.5 Text-Only Performance
Table 6 compares jina-vlm against the backbone Qwen3-1.7B on text-only benchmarks: MMLU (mmlu), MMLU-Pro (mmlupro), GSM-8K (gsm8k), ARC-C (arc), and HellaSwag (hellaswag). jina-vlm trails slightly on knowledge-intensive tasks (MMLU, GSM-8K) but matches or exceeds the backbone on ARC-C and HellaSwag, indicating minimal degradation from multimodal training.
@@ 6 Conclusion
We presented jina-vlm, a 2.4B vision-language model that achieves state-of-the-art multilingual visual question answering among open models at comparable scale. Our results demonstrate that small VLMs can attain strong cross-lingual visual understanding through careful architectural and training choices: attention-based token pooling reduces visual tokens by 4 while preserving spatial information, and incorporating text-only data during multimodal training mitigates the catastrophic forgetting typically observed in vision-adapted language models. On standard English VQA benchmarks, jina-vlm matches or exceeds models of similar size, suggesting that multilingual capabilities need not come at the cost of general performance.
The current approach has limitations. Multi-tile processing introduces computational overhead that scales with image resolution, and we have not emphasized safety-critical training or alignment. Future work could explore more efficient resolution handling and investigate whether our multilingual training recipe transfers to larger model scales.

### v2

@@ Abstract
We present jina-vlm, a 2.4B parameter vision-language model that achieves state-of-the-art multilingual visual question answering among open 2B-scale VLMs. The model couples a SigLIP2 vision encoder with a Qwen3 language backbone through an attention-pooling connector that enables token-efficient processing of arbitrary-resolution images. The model achieves leading results on standard VQA benchmarks and multilingual evaluations while preserving competitive text-only performance. Model weights and code are publicly released at https://huggingface.co/jinaai/jina-vlm.
@@ 1 Introduction
This work introduces jina-vlm, a 2.4B parameter VLM that addresses both challenges. The model aligns a SigLIP2-So400M/14-384 vision encoder (siglip2) with Qwen3-1.7B-Base (qwen3) through an attention-pooling connector, trained with a two-stage pipeline that explicitly incorporates multilingual data. Among open 2B-scale VLMs, jina-vlm achieves state-of-the-art performance on multilingual multimodal benchmarks including MMMB and Multilingual MMBench, demonstrating that small models can excel at cross-lingual visual understanding without sacrificing general capabilities. On standard English benchmarks spanning diagrams, charts, documents, and OCR, jina-vlm achieves the highest average score (72.3) across eight VQA benchmarks among 2B-scale VLMs. These results are enabled by two technical contributions: an efficient arbitrary-resolution pipeline that combines overlapping tiling with attention-based token pooling to reduce visual token count by 4, and a training recipe that incorporates text-only data to preserve the language understanding performance of the backbone LLM.
@@ 2 Related Work
Efficient resolution-agnostic image processing. Standard ViTs process fixed-resolution images, requiring resizing that discards fine-grained detail. Since visual token count scales with resolution and Transformer computation scales quadratically with sequence length, naive high-resolution processing is prohibitive. Several solutions exist: molmo tile images with overlap; qwen_2_vl introduce Naive Dynamic Resolution with Multimodal Rotary Position Embedding (rope; 2drope); ovis_2_5 use native-resolution ViTs (navit). Orthogonally, images often contain low-information regions (e.g., sky backgrounds), making visual tokens highly redundant. Token compression methods address this (fastv; prumerge; visionzip; pyramiddrop). internvl_1_5 develop Dynamic High-Resolution Tiling, and nvila propose scale-then-compress strategies. Recent work on training-free token budgeting, such as HERO (hero), demonstrates that inference-time pruning can achieve significant speedups while preserving accuracy; our approach differs by learning compact representations during training rather than dropping tokens at inference.
Vision-language connectors. The connector bridging vision encoders and language models significantly impacts both efficiency and performance. BLIP-2 (blip2) introduces Q-Former, a learnable query-based transformer that extracts fixed-length representations from visual features, reducing the number of tokens fed to the LLM. Flamingo (flamingo) uses a Perceiver Resampler with cross-attention to compress visual tokens. Our attention-pooling connector shares the goal of token reduction but operates differently: rather than learning a fixed set of queries, we apply local 22 attention pooling that preserves spatial structure while achieving 4 compression, which we found more effective for tasks requiring fine-grained spatial understanding.
@@ 3.2 Language Decoder
3.3 Efficiency Analysis
Table 1 quantifies the computational benefits of attention pooling. With the default 12-tile configuration (plus thumbnail), the unpooled baseline would produce 9,477 visual tokens per image, while our 22 pooling reduces this to 2,366 tokens. Since the ViT processes each tile identically regardless of pooling, the savings apply exclusively to the LLM: we observe a 3.9 reduction in prefill FLOPs and a 4 reduction in KV-cache memory. The overall FLOPs reduction is 2.3 when including the shared ViT cost.
| Metric | No Pooling | With Pooling | Reduction |
| Visual tokens | 9,477 | 2,366 | 4.0 |
| LLM prefill FLOPs | 27.2 TFLOPs | 6.9 TFLOPs | 3.9 |
| KV-cache memory | 2.12 GB | 0.53 GB | 4.0 |
@@ 5.1 General VQA Tasks
Results for models other than jina-vlm are from their respective papers, internvl_3_5; internvl_3; qwen_2_vl, except those marked with * which were computed using VLMEvalKit. All scores represent accuracy (%) except OCRBench which uses a 0–1000 scale; for overall average computation, OCRBench scores are divided by 10 to align with the 0–100 scale of other benchmarks.
@@ 5.2 Document and Real-World Understanding
Results for models other than jina-vlm are from their respective papers, internvl_3_5; internvl_3; qwen_2_vl, except those marked with * which are computed using VLMEvalKit. All scores represent accuracy (%) except MME which uses a 0–2800 scale; for overall average computation, MME scores are divided by 28 to align with the 0–100 scale of other benchmarks.
@@ 5.5 Text-Only Performance
Table 7 compares jina-vlm against the backbone Qwen3-1.7B on text-only benchmarks: MMLU (mmlu), MMLU-Pro (mmlupro), GSM-8K (gsm8k), ARC-C (arc), and HellaSwag (hellaswag). Results show mixed preservation of text-only capabilities: jina-vlm matches or exceeds the backbone on commonsense reasoning (ARC-C, HellaSwag) and retains most performance on MMLU and GSM-8K. However, MMLU-Pro shows substantial degradation (46.4 30.3), likely because this benchmark emphasizes extended multi-step reasoning that conflicts with our instruction-tuning toward concise visual responses. This suggests a trade-off between optimizing for multimodal tasks and preserving complex text-only reasoning, which future work could address through more balanced data mixtures or curriculum scheduling.
@@ 6 Conclusion
We presented jina-vlm, a 2.4B vision-language model that achieves state-of-the-art multilingual visual question answering among open 2B-scale VLMs. Our results demonstrate that small VLMs can attain strong cross-lingual visual understanding through careful architectural and training choices: attention-based token pooling reduces visual tokens by 4 while preserving spatial information, and incorporating text-only data during multimodal training mitigates the catastrophic forgetting typically observed in vision-adapted language models. On standard English VQA benchmarks, jina-vlm achieves leading results, demonstrating that multilingual capabilities need not come at the cost of general performance.
The current approach has limitations. Multi-tile processing introduces computational overhead that scales with image resolution, and tiling can fragment global spatial context, potentially impairing performance on tasks requiring holistic scene understanding such as object counting or precise spatial reasoning across tile boundaries. While the global thumbnail partially mitigates this, native-resolution approaches (navit) may be better suited for such tasks. We have not emphasized safety-critical training or alignment, and multi-image reasoning remains weak due to limited training data in this regime. Future work could explore more efficient resolution handling, targeted improvements for counting and spatial tasks, and investigate whether our multilingual training recipe transfers to larger model scales.